Find the relative extreme values of the function f (x) = x3 − x2 − 6x + 5.


1
Expert's answer
2021-05-26T15:46:28-0400



Given

f(x)=x3-x2-6x+5


We will find extreme

First of find 1st derivative of f(x)


f'(x)=3x2-2x-6


Put f'(x) equal to zero for extremize.

f'(x)=0

3x2-2x-6=0


Now find value of x.

"x=\frac{[2± \sqrt{4-4.3(-6)}]}{2.3}\\

x=\frac{(1±\sqrt{\smash[b]{19}}) }{3}"





this values of x give us extreme value of f(x)


Now find second derivative of f(x) for find maxima or minima.


f''(x)= 6x-2


Now check 

"x=\frac{(1+\sqrt{\smash[b]{19}}) }{3}"


is give maxima or minima.

So put this value in f''(x)

"=6(\frac{(1+\sqrt{\smash[b]{19}}) }{3})-2"


=8.717

It is positive value .it mean this value of x give us minima of f(x).

Put this value of x in f(x).

"f(x)=(\frac{1+\sqrt{\smash[b]{19}})}{3})^3-(\frac{1+\sqrt{\smash[b]{19}})}{3})^2-6(\frac{1+\sqrt{\smash[b]{19}})}{3})+5"


=9.59802825019

This is minima value.


Now check

"x=\frac{(1-\sqrt{\smash[b]{19}}) }{3}"

 

is give maxima or minima.

So put this value in f''(x)

"=6(\frac{(1-\sqrt{\smash[b]{19}}) }{3})-2"


=-8.71779788708

It is negative value .it mean this value of x give us maxima of f(x).

Put this value of x in f(x).

"f(x)=(\frac{1-\sqrt{\smash[b]{19}})}{3})^3-(\frac{1-\sqrt{\smash[b]{19}})}{3})^2-6(\frac{1-\sqrt{\smash[b]{19}})}{3})+5"


=11.7984052203

This is maxima value.